there are 3 times as many pears as oranges. If a group of children receives 5 organges each, there will be no oranges left over. If the same group of children receives 8 pears each, there will be 21 pears left over. How many children and organges are there?
Answer by venugopalramana(3286)

there are 3 times as many pears as oranges. If a group of children receives 5 organges each, there will be no oranges left over. If the same group of children receives 8 pears each, there will be 21 pears left over. How many children and organges are there?
LET ORANGES=X
PEARS ARE 3 TIMES =3X
LET CHILDREN=Y
IF WE GIVE 5 ORANGES PER CHILD WE NEED 5Y ORANGES.SO THIS =X
X=5Y........................I
IF WE GIVE 8 PEARS PER CHILD WE NEED 8Y PERARS...THERE ARE STILL 21 PEARS LEFT OUT AFTER THIS THAT IS
8Y+21=3X....................II
SUBSTITUTING FOR X FROM EQN.I..
8Y+21=3*5Y
15Y-8Y=21
7Y=21
Y=3
X==5Y=5*3=15..
SO CHILDREN =3
ORANGES=15
PEARS=3*15=45